OpenStudy (anonymous):

What is the value of i ^5

OpenStudy (anonymous):

well what is i?

OpenStudy (anonymous):

-i

OpenStudy (anonymous):

sry i

OpenStudy (anonymous):

yeah, assuming you mean i as the square root of -1, then i^0 = 1 (any # to the power of 0 is one) i ^1 = i i^2 = -1 (the square root goes away) i^3 = -i (since the left hand side is equivalent to i^2 * i = -i ) From then on the values loop, so i^4 = 1, i^5 = i, etc. You can also make them up from the values of previous terms to verify, i.e. i^5 = i^3 * i^2 = -1*-i = i

OpenStudy (anonymous):

Is it Diff E problem?